what is the inverse of f(x)=5-3x?
1
Expert's answer
2014-07-10T02:05:26-0400
Solve y=5-3x for "x=": y-5=-3x --> x=(5-y)/3. Then the inverse of f(x)=5-3x is f(x)=(5-x)/3.